Darla owns a dress shop called Darla's Darling Dresses. During the past year, Darla sold some assets to upgrade her facility. She sold racks and shelving units for $600 cash. In addition to the cash, the buyer gave two mannequins worth a total of $200 to Darla. The racks and shelving units had an original cost of $2,500 and had accumulated depreciation for tax purposes of $2,200.

Darla's amount realized on the sale is $:

and the adjusted basis in the assets sold is $:

, producing a realized on the sale of $:
Answer #1

Darla's amount realized on the sale is $800 and the adjusted basis in the assets sold is $300, producing a realized gain on the sale of $500

Amount realized = cash received + FMV of other property + buyer’s assumption of seller’s liabilities – seller’s expenses

Amount realized = 600 + 200 + 0 -0 = $800

Adjusted basis = initial basis – cost recovery deductions

Adjusted basis = 2500-2200 = $300

Gain or loss realized = amount realized – adjusted basis = 800-300 = $500